# Provably Efficient RLHF Pipeline: A Unified View from Contextual Bandits

### v1

@@ 3 Problem Setup
Through direct calculation, we obtain $\kappa\leq 2+\exp(2BL)+\exp(-2BL)$. Thus, $\kappa$ may be exceedingly large with an exponential dependence on the magnitude of features and parameters, as justified in Appendix G.2.
The RLHF pipeline can generally be divided into two distinct stages: the (post-)training stage and the deployment stage, each with different objectives. In the training stage, the learner aims to develop a high-performing model with minimal samples and computational resources. In the deployment stage, the trained model is used to generate good responses to user inputs, while simultaneously gathering informative feedback for further refinement. We provide a detailed discussion of both stages in the following sections.
4 Training Stage
In the training stage, the learner aims to acquire a high-performing model with minimal number of samples and computational resources. To effectively evaluate the model’s performance, we define the following performance measure.
Performance Measure.
For any policy $\pi$, we define the sub-optimality gap compared to the optimal policy as follows:
$\operatorname*{\mathsf{SubOpt}}(\pi)=\mathbb{E}_{x\sim\rho}\left[r\left(x,\pi^{*}(x)\right)-r(x,\pi(x)\right],$ (3)
where $\rho$ is the context distribution, $\pi^{*}=\arg\max_{a}r(x,a)$ is the optimal policy under the true reward $r$. The goal of the learner is to minimize the performance gap $\operatorname*{\mathsf{SubOpt}}(\pi)$.
In the training stage, two data collection strategies are commonly used: passive and active. In passive setting, the learner is given a fixed set of samples. Conversely, in active setting, the learner can actively select queries for human feedback. We discuss both strategies below.
4.1 Passive Data Collection
In the passive data collection setting, we are given a fixed dataset $\left\{\left(x_{t},a_{t},a_{t}^{\prime},y_{t}\right)\right\}_{t=1}^{T}$ where $x_{t}\in\mathcal{X}$ is the context, $a_{t},a_{t}^{\prime}\in\mathcal{A}$ are the actions, and $y_{t}\in\{0,1\}$ is the feedback. Since the parameter $\theta^{*}$ is unknown, we need to estimate it from the available data. The most common method for this estimation is maximum likelihood estimation (MLE). First, let $z=\phi(x,a)-\phi(x,a^{\prime})$ be the feature difference, then we define the negative log-likelihood $\ell_{t}(\theta)$ for sample $t$ as
$\ell_{t}(\theta)=-y_{t}\log(\sigma(z_{t}^{\top}\theta))-\left(1-y_{t}\right)\log(1-\sigma(z_{t}^{\top}\theta)).$ (4)
The MLE estimator $\widehat{\theta}_{T+1}$ is then defined as
$\widehat{\theta}_{T+1}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\mathcal{L}_{T+1}(\theta)\triangleq\sum_{t=1}^{T}\ell_{t}(\theta).$ (5)
Several works (Zhu et al., 2023; Ji et al., 2024) have analyzed the statistical properties of MLE and showed that for any $\delta\in(0,1)$, with probability at least $1-\delta$, the MLE estimator $\widehat{\theta}_{T+1}$ converges to the true parameter $\theta^{*}$ by
$\lVert\widehat{\theta}_{T+1}-\theta^{*}\rVert_{V_{T+1}}\leq\widehat{\beta}\triangleq\widetilde{\mathcal{O}}\big{(}\kappa\sqrt{d}\big{)},$ (6)
where $V_{t}\triangleq\sum_{i=1}^{t-1}z_{i}z_{i}^{\top}+\lambda I$ is the feature covariance matrix, $\lambda$ is the regularization parameter.
For computational complexity, the MLE estimator does not have a closed-form solution, requiring iterative optimization techniques, such as gradient descent, to obtain an $\varepsilon$-accurate solution. As discussed in Faury et al. (2022), optimizing the MLE typically requires $\mathcal{O}(\log(1/\varepsilon))$ iterations to reach the desired accuracy. Since the loss function is defined over all data samples, each iteration involves $\mathcal{O}(T)$ gradient computations. Given that $\varepsilon$ is typically set to $1/T$ to prevent it from dominating the overall estimation error, the total computational complexity is $\mathcal{O}(T\log T)$.
Although the MLE estimator is a commonly used approach, we propose a novel algorithm that enhances both computational efficiency and sample efficiency compared to MLE, drawing inspiration from recent advancements in logistic bandits (Faury et al., 2022; Zhang and Sugiyama, 2023) and multinomial logit MDPs (Li et al., 2024). First, we define the gradient $g_{t}(\theta)$ and Hessian $H_{t}(\theta)$ of the loss function $\ell_{t}(\theta)$ as
$g_{t}(\theta)=\left(\sigma\left(z_{t}^{\top}\theta\right)-y_{t}\right)z_{t},~{}~{}H_{t}(\theta)=\dot{\sigma}\left(z_{t}^{\top}\theta\right)z_{t}z_{t}^{\top}.$ (7)
Implicit OMD.
To improve the computational efficiency, Faury et al. (2022) observed that the cumulative past log-loss is strongly convex and can therefore be well approximated by a quadratic function. Building on this observation, they proposed the following update rule:
where ¯$\bar{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}\left(\bar{\theta}_{i+1}\right)+\lambda I$ is the local norm, and $\eta$ is the step size. The optimization problem can be decomposed in two terms. The first term is the instantaneous log-loss $\ell_{t}(\theta)$, which accounts for the information of the current sample. The second consists of a quadratic proxy for the past losses constructed through the sequence $\{\bar{\theta}_{i}\}_{i\leq t}$. A key component is the design of the local norm ¯$\bar{\mathcal{H}}_{t}$, which approximates the Hessian matrix by $H_{i}\left(\bar{\theta}_{i+1}\right)$ at a lookahead point $\bar{\theta}_{i+1}$. Such a Hessian matrix effectively captures local information and is crucial for ensuring statistical efficiency.
Although this formulation may appear computationally more efficient than MLE due to its one-pass data processing, the optimization problem in Eq. (8) still lacks a closed-form solution. Consequently, it still requires $\mathcal{O}(\log(1/\varepsilon))$ iterations per sample to achieve an $\varepsilon$-accurate solution, leading to the same $\mathcal{O}(T\log T)$ computational complexity as MLE.
Standard OMD.
To enhance computational efficiency, we revisit the formulation in Eq. (8) and observe that it essentially corresponds to the implicit OMD framework (Kivinen and Warmuth, 1997). Consequently, a natural alternative is to replace this formulation with the standard OMD framework, which permits a closed-form solution and thus eliminates the need for iterative optimization. However, the standard OMD minimizes a first-order approximation of the loss function, which sacrifices several key properties compared to its implicit counterpart, as demonstrated by Campolongo and Orabona (2020). Specifically, the standard OMD formulation updates using $\partial\ell_{t}(\theta_{t})$, whereas the implicit OMD updates the algorithm approximately with the subsequent sub-gradient, $\partial\ell_{t}(\theta_{t+1})$. This distinction results in a notable gap in the convergence rates of the two methods. To this end, we propose to approximate the current loss $\ell_{t}(\theta)$ using a second-order Taylor expansion, drawing inspiration from Zhang and Sugiyama (2023). We define the second-order approximation of the original loss function $\ell_{t}(\theta)$ at the current estimator $\widetilde{\theta}_{t}$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+g_{t}(\widetilde{\theta}_{t})^{\top}(\theta-\widetilde{\theta}_{t})+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}$. Then, we replace the current loss $\ell_{t}(\theta)$ in Eq. (8) with the approximation ~$\widetilde{\ell}_{t}(\theta)$, leading to the following update rule:
where $\eta$ is the step size and ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ is the local norm with $\mathcal{H}_{t}\triangleq\mathcal{H}_{t}(\widetilde{\theta}_{i+1})=\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\lambda I$.
The optimization problem in Eq. (9) can be solved with a projected gradient step with the following equivalent formulation:
$\widetilde{\theta}_{t+1}^{\prime}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t}),\quad\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\widetilde{\theta}_{t+1}^{\prime}\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}.$
Thus, the estimator $\widetilde{\theta}_{t+1}$ can be computed in an $\mathcal{O}(1)$ time complexity, leading to a total time complexity of $\mathcal{O}(T)$ for the entire procedure. This represents a significant improvement over the $\mathcal{O}(T\log T)$ complexity of the MLE estimator in Eq. (5), particularly considering that the number of samples $T$ is typically large in the training of LLMs. Furthermore, since the estimator processes the samples in a one-pass manner, it mitigates the memory burden associated with computing the gradient of the full dataset at each iteration. This makes it more suitable for large-scale applications, where memory and time efficiency are crucial.
It is important to note that the update rule in Eq. (9) represents a special case of online mirror descent, specifically, it can be derived from the following general form:
where $\psi_{t}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$ is the regularizer and $\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})=\psi_{t}(\theta)-\psi_{t}(\widetilde{\theta}_{t})-\langle\nabla\psi_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle$ is the Bregman divergence. Therefore, we can analyze the estimation error using the online mirror descent framework (Orabona, 2019). Specifically, we have the following lemma.
Algorithm 1 Training Stage: Passive Data Collection
0:  Dataset $\left\{\left(x_{t},a_{t},a_{t}^{\prime},y_{t}\right)\right\}_{t=1}^{T}$, regularization $\lambda>0$
1:  Initialize $\widetilde{\theta}_{1}=0$, $\mathcal{H}_{1}=\lambda I_{d}$.
@@ 4.1 Passive Data Collection
3:     Define $g_{t}(\theta)$ and $H_{t}(\theta)$ as in Eq. (7)
4:     Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
5:     Compute $\widetilde{\theta}_{t+1}$ by Eq. (9)
6:     Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
@@
8:  Compute $\widetilde{J}_{T}(\pi)$ by Eq. (11)
8:  Compute final policy $\pi_{T}=\operatorname*{arg\,max}_{\pi}\widetilde{J}_{T}(\pi)$
Lemma 1.
Let $\delta\in(0,1]$, set $\eta=(1/2)\log 2+\left(BL+1\right)$ and $\lambda=84\sqrt{2}\eta(dL^{2}+BL^{3})$, define the confidence set as
$\mathcal{C}_{t}=\left\{\theta\in\Theta\mid\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\mathcal{H}_{t}}\leq\widetilde{\beta}_{t}\triangleq\mathcal{O}\big{(}\sqrt{d}(\log(t/\delta))^{2}\big{)}\right\}.$
Then, we have $\operatorname{Pr}\left[\forall t\geqslant 1,\theta^{*}\in\mathcal{C}_{t}(\delta)\right]\geqslant 1-\delta$.
Intuitively, the confidence set shrinks as the number of samples $t$ increases, owing to the growing local norm $\mathcal{H}_{t}$. The local norm $\mathcal{H}_{t}$ captures the directions explored by the samples. The more we explore in a particular direction, the more accurate our estimation becomes along that direction.
By the definition of $\kappa$ in Definition 2, it holds that $\mathcal{H}_{t}\succeq(1/\kappa)V_{t}$, Lemma 1 implies $\forall t\in[T+1]$,
$\lVert\widetilde{\theta}_{t}-\theta^{*}\rVert_{V_{t}}\leq\sqrt{\kappa}\lVert\widetilde{\theta}_{t}-\theta^{*}\rVert_{\mathcal{H}_{t}}\leq\widetilde{\mathcal{O}}\big{(}\sqrt{\kappa d}\big{)}.$ (10)
Compared with the confidence set in Eq. (6), our confidence set strictly improves the estimation error by a factor of $\sqrt{\kappa}$. As highlighted in Remark 2, the non-linearity coefficient $\kappa$ may be exceedingly large during the training of LLMs, which makes our estimator statistically more efficient.
Pessimistic Policy.
Based on the estimator in Eq. (9), we adopt the “pessimism in the face of uncertainty” principle and define the pessimistic value function as
$\widetilde{J}_{T}(\pi)=\min_{\theta\in\mathcal{C}_{T}}\mathbb{E}_{x\sim\rho}\left[(\phi(x,\pi(x))-v)^{\top}\theta\right]=(\bar{\phi}(\pi)-v)^{\top}\widetilde{\theta}_{T}-\widetilde{\beta}_{T}\lVert\bar{\phi}(\pi)-v\rVert_{\mathcal{H}_{T}^{-1}},$ (11)
where $\bar{\phi}(\pi)=\mathbb{E}_{x\sim\rho}[\phi(x,\pi(x))]$ is the feature expectation of policy $\pi$, and $v$ is the baseline feature vector. While $v$ does not affect the induced policy, it may help reduce the sub-optimality gap (Zhu et al., 2023). The final policy is given by $\pi=\operatorname*{arg\,min}_{\pi}\widetilde{J}(\pi)$. The procedure is present in Algorithm 1. We show it enjoys the following guarantee.
Theorem 1.
Let $\delta\in(0,1]$, set the parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 1 ensures the following sub-optimality gap:
${\operatorname*{\mathsf{SubOpt}}}(\pi_{T})\leq\widetilde{\mathcal{O}}\left(\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}\right).$
Remark 3.
The term $\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{\mathcal{H}_{T}^{-1}}$ is referred to as the “concentratability coefficient”, a widely used concept in offline learning (Jin et al., 2021; Zhu et al., 2023). It measures the distribution shift between the optimal policy and the dataset distribution. A smaller concentratability coefficient value indicates that the dataset better covers the data distribution under the optimal policy, resulting in a smaller sub-optimality gap.
Comparison with Zhu et al. (2023).
Our result improves upon their work in both statistical and computational efficiency. First, similar to Eq. (10), Theorem 1 implies an $\widetilde{\mathcal{O}}\big{(}\sqrt{\kappa d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left(\phi(x,\pi^{*}(x))-v\right)\big{\rVert}_{V_{T}^{-1}}\big{)}$ sub-optimality gap, improving upon theirs by a $\sqrt{\kappa}$ factor. Moreover, our estimator has an $\mathcal{O}(T)$ time complexity, more efficient than their $\mathcal{O}(T\log T)$ complexity.
4.2 Active Data Collection
In this part, we consider the active data collection setting, where the learner can actively select tuple $\{x_{t},a_{t},a^{\prime}_{t}\}_{t=1}^{T}$ to present to the human for feedback $\{y_{t}\}_{t=1}^{T}$.
Estimator Construction.
For each round $t\in[T]$, given $\left\{\left(x_{i},a_{i},a_{i}^{\prime},y_{i}\right)\right\}_{i=1}^{t-1}$, we define the loss function $\ell_{t}(\theta)$ as in Eq. (4), then the most standard approach is to estimate the parameter $\theta^{*}$ by MLE, which is given by $\widehat{\theta}_{t}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{t-1}\ell_{i}(\theta)$. However, as noted before, the MLE estimator leads to a time complexity of $\mathcal{O}(t\log t)$ per round. Moreover, unlike the passive setting where the estimator is constructed only once, in the active setting, we need to construct the estimator in each round. This is prohibitively expensive for large-scale applications.
To improve the computational efficiency, we can leverage the OMD framework as in the passive data collection setting. Specifically, we construct the estimator $\widetilde{\theta}_{t}$ as in Eq. (9) by
$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}}.$ (12)
Algorithm 2 Training Stage: Active Data Collection
0:  Regularization $\lambda>0$, failure probability $\delta\in(0,1]$
1:  Initialize $\widetilde{\theta}_{1}=0$, $\mathcal{H}_{1}=\lambda I_{d}$.
2:  for $t=1,2,\ldots,T$ do
3:     Choose $(x_{t},a_{t},a_{t}^{\prime})$ by Eq. (13), observe feedback $y_{t}$
4:     Define $g_{t}(\theta)$ and $H_{t}(\theta)$ as in Eq. (7)
5:     Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
6:     Compute $\widetilde{\theta}_{t+1}$ by Eq. (12)
7:     Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
8:  end for
9:  Construct $\widetilde{r}_{T}(x,a)=\frac{1}{T}\sum_{t=1}^{T}\phi(x,a)^{\top}\widetilde{\theta}_{t},\forall x,a$
9:  Compute policy $\pi_{T}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T}(x,a)$.
Query Selection.
Given the estimator $\widetilde{\theta}_{t}$, we choose the query with the maximum uncertainty inspired by (Das et al., 2024). Specifically, we choose the tuple by

### v2

@@ 3 Problem Setup
The quantity $\kappa$ captures the learning complexity that satisfies $\kappa\leq 2+\exp(2BL)+\exp(-2BL)$. It may be exceedingly large with exponential dependence on the scale of features and parameters.
4 Our Framework
In this section, we first introduce the general framework for online RLHF. We then present our one-pass reward modeling method. Finally, we show the theoretical guarantee of our method.
4.1 General Framework for Online RLHF
The general process of online RLHF involves iteratively collecting data and updating the model based on the collected data. At iteration $t$, the process can be formulated as:
  1. (i)
    New data collection: Sample a prompt $x_{t}$ and two responses $a_{t}$ and $a_{t}^{\prime}$, query the oracle to obtain the preference label $y_{t}\in\{0,1\}$, expand the dataset $\mathcal{D}_{t+1}=\mathcal{D}_{t}\cup\{(x_{t},a_{t},a_{t}^{\prime},y_{t})\}$.
  2. (ii)
    Reward modeling: Train a reward model $r_{t+1}$ using the historical dataset $\mathcal{D}_{t+1}$.
  3. (iii)
    Policy optimization (Optional): Update the policy $\pi_{t+1}$ using the reward model $r_{t+1}$.
A key challenge in online RLHF is that the reward model needs to be trained on the entire historical dataset at each iteration, which is computationally expensive. Specifically, let $z_{t}=\phi(x_{t},a_{t})-\phi(x_{t},a_{t}^{\prime})$ be the feature difference, given the historical dataset $\mathcal{D}_{t+1}=\left\{(x_{i},a_{i},a_{i}^{\prime},y_{i})\right\}_{i=1}^{t}$, the reward model is estimated via maximum likelihood estimation as
$\widehat{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\sum_{i=1}^{t}\ell_{i}(\theta),\text{where }\ell_{t}(\theta)=-y_{t}\log(\sigma(z_{t}^{\top}\theta))-\left(1-y_{t}\right)\log(1-\sigma(z_{t}^{\top}\theta)).$ (3)
However, Eq. (3) does not admit a closed-form solution, requiring iterative optimization techniques, such as gradient descent, to achieve an $\varepsilon$-accurate estimate. As discussed by Faury et al. (2022), obtaining such accuracy with MLE typically requires $\mathcal{O}(\log(1/\varepsilon))$ optimization steps. Since the loss function is defined over the entire historical dataset, each iteration incurs a computational cost of $\mathcal{O}(t)$ gradient evaluations. In practice, $\varepsilon$ is often set to $1/t$ to ensure that the optimization error does not dominate the overall estimation error. As a result, the total computational complexity at iteration $t$ becomes $\mathcal{O}(t\log t)$, a cost that is prohibitive for long-term online RLHF applications.
4.2 One-pass Reward Modeling
Drawing inspiration from recent advancements in logistic bandits (Faury et al., 2022; Zhang and Sugiyama, 2023) and multinomial logit MDPs (Li et al., 2024), we propose a novel one-pass reward modeling method that reduces the complexity to constant time per iteration. First, define the gradient $g_{t}(\theta)$ and Hessian $H_{t}(\theta)$ of loss $\ell_{t}(\theta)$ as $g_{t}(\theta)=(\sigma(z_{t}^{\top}\theta)-y_{t})z_{t}$ and $H_{t}(\theta)=\dot{\sigma}(z_{t}^{\top}\theta)z_{t}z_{t}^{\top}$.
Implicit OMD.  To improve the computational efficiency, Faury et al. (2022) observed that the cumulative past log-loss is strongly convex and can therefore be well approximated by a quadratic function. Building on this observation, they proposed the following update rule:
where ¯$\bar{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}\left(\bar{\theta}_{i+1}\right)+\lambda I$ is the local norm, and $\eta$ is the step size. The optimization problem can be decomposed into two terms. The first term is the instantaneous log-loss $\ell_{t}(\theta)$, which accounts for the information of the current sample. The second consists of a quadratic proxy for the past losses constructed through the sequence $\{\bar{\theta}_{i}\}_{i\leq t}$. A key component is the design of the local norm ¯$\bar{\mathcal{H}}_{t}$, which approximates the Hessian matrix by $H_{i}\left(\bar{\theta}_{i+1}\right)$ at a lookahead point $\bar{\theta}_{i+1}$. Such a Hessian matrix effectively captures local information and is crucial for ensuring statistical efficiency.
The update rule in Eq. (4) benefits from a one-pass data processing property, which eliminates the need to store the entire historical dataset. However, the optimization problem in Eq. (4) still does not have a closed-form solution. But since the loss is defined only on the current sample, it requires only $\mathcal{O}(1)$ gradient computations per step, leading to a total computational complexity of $\mathcal{O}(\log t)$ at iteration $t$. This represents a significant improvement over the $\mathcal{O}(t\log t)$ complexity of the MLE estimator in Eq. (3). Nevertheless, the computational complexity of the implicit OMD is still increasing with the number of iterations, which motivates us to design a constant-time method.
Standard OMD.  To enhance computational efficiency, a natural alternative is to replace this formulation with the standard OMD framework, which permits a closed-form solution and thus eliminates the need for iterative optimization. However, the standard OMD minimizes a first-order approximation of the loss function, which sacrifices several key properties compared to its implicit counterpart, as demonstrated by Campolongo and Orabona (2020). Specifically, the standard OMD formulation updates using $g_{t}(\theta_{t})$, whereas the implicit OMD updates the algorithm approximately with the subsequent sub-gradient, $g_{t}(\theta_{t+1})$. This distinction results in a notable gap in the convergence rates of the two methods. To this end, we propose to approximate the current loss $\ell_{t}(\theta)$ using a second-order Taylor expansion, drawing inspiration from Zhang and Sugiyama (2023). Define the second-order approximation of $\ell_{t}(\theta)$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+g_{t}(\widetilde{\theta}_{t})^{\top}(\theta-\widetilde{\theta}_{t})+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}$. Then, we replace the loss $\ell_{t}(\theta)$ in Eq. (4) with the approximation ~$\widetilde{\ell}_{t}(\theta)$, leading to the update rule:
where $\eta$ is the step size and ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ is the local norm with $\mathcal{H}_{t}\triangleq\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\lambda I$. Eq. (5) can be solved with a projected gradient step with the following equivalent form:
$\widetilde{\theta}_{t+1}^{\prime}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t}),~{}~{}~{}\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\widetilde{\theta}_{t+1}^{\prime}\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}.$
Thus, the estimator $\widetilde{\theta}_{t+1}$ provides a closed-form solution, leading to a $\mathcal{O}(1)$ computational complexity per iteration. Since the estimator processes the samples in a one-pass manner, it mitigates the memory burden associated with computing the gradient of the full dataset. These properties make the method particularly suitable for edge devices, where both memory and computational resources are severely constrained. The detailed process of our proposed method is presented in Algorithm 1.
Algorithm 1 One-Pass Reward Modeling
0:  Preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$
1:  Define the loss function $\ell_{t}(\theta)$ as Eq. (3)
2:  Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
3:  Compute $\widetilde{\theta}_{t+1}^{\prime}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$
4:  Compute $\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\lVert\theta-\widetilde{\theta}_{t+1}^{\prime}\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$
5:  Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
5:  $\widetilde{\theta}_{t+1}$
Algorithm 2 Passive Data Collection
0:  Regularization $\lambda$, step size $\eta$
1:  Initialize $\widetilde{\theta}_{1}=\bm{0}$ and ~$\widetilde{\mathcal{H}}_{1}=\lambda I$
2:  for $t=1,2,\ldots,T$ do
3:     Observe preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$
[…]
5:  end for
6:  Construct $\widetilde{J}_{T+1}(\pi)$ as in Eq. (6)
6:  $\pi_{T+1}=\operatorname*{arg\,max}_{\pi\in\Pi}\widetilde{J}_{T+1}(\pi)$
4.3 Theoretical Guarantee
Note that the update rule in Eq. (5) is a special case of online mirror descent, specifically:
where $\psi_{t}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$ is the regularizer and $\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})=\psi_{t}(\theta)-\psi_{t}(\widetilde{\theta}_{t})-\langle\nabla\psi_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle$ is Bregman divergence. Based on the analysis of online mirror descent, we have the following lemma.
Lemma 1.
Let $\delta\in(0,1]$, set $\eta=(1/2)\log 2+\left(BL+1\right)$ and $\lambda=84\sqrt{2}\eta(dL^{2}+BL^{3})$, define $\mathcal{C}_{t}=\{\theta\in\Theta\mid\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\beta}_{t}\triangleq\mathcal{O}\big{(}\sqrt{d}(\log(t/\delta))^{2}\big{)}\}$. Then, we have $\operatorname{Pr}\left[\forall t\geqslant 1,\theta^{*}\in\mathcal{C}_{t}\right]\geqslant 1-\delta$.
Comparison with MLE.  For the MLE estimator in Eq. (3), prior works (Zhu et al., 2023; Das et al., 2024; Ji et al., 2024) have shown $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\widetilde{\mathcal{O}}(\kappa\sqrt{d})$, where $V_{t}=\sum_{i=1}^{t-1}z_{i}z_{i}^{\top}+\lambda I$. By the definition of $\mathcal{H}_{t}$, it holds that $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Lemma 1 implies $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\sqrt{\kappa}\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\mathcal{O}}(\sqrt{\kappa d})$. This result shows that Lemma 1 improves upon previous bounds by at least a factor of $\sqrt{\kappa}$.
5 Applications in Three Online RLHF Scenarios
In this section, we apply our framework to three distinct RLHF scenarios, including online RLHF with passive data collection, active data collection, and deployment-time adaptation.
(a) Passive Data Collection
(b) Active Data Collection
(c) Deployment-Time Adaptation
[…]
5.1 Online RLHF with Passive Data Collection
We first consider the passive data collection setting, where the algorithm can not control the data collection process. At each iteration, the learner obtains $(x_{t},a_{t},a_{t}^{\prime},y_{t})$ and updates by Eq. (5). We adopt the “pessimism in the face of uncertainty” principle and define the value function $\widetilde{J}_{t+1}(\pi)$ as
$\widetilde{J}_{T+1}(\pi)=(\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi(x))\right])^{\top}\widetilde{\theta}_{T+1}-\widetilde{\beta}_{T+1}\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi(x))\right]\rVert_{\mathcal{H}_{T+1}^{-1}}.$ (6)
where $\rho$ is the context distribution. The policy $\pi_{T+1}$ is selected as $\pi_{T+1}=\operatorname*{arg\,max}_{\pi\in\Pi}\widetilde{J}_{T+1}(\pi)$. The detailed procedure is present in Algorithm 2, and we show it enjoys the following guarantee:
Theorem 1.
Set parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 2 ensures
${\operatorname*{\mathsf{SubOpt}}}(\pi_{T+1})=\mathbb{E}_{x\sim\rho}\left[r(x,\pi^{*}(x))-r(x,\pi_{T+1}(x))\right]\leq\widetilde{\mathcal{O}}\left(\sqrt{d}\cdot\big{\lVert}\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\big{\rVert}_{\mathcal{H}_{T+1}^{-1}}\right),$
where $\rho$ is the context distribution and $\pi^{*}$ is the optimal policy.
@@ 5.1 Online RLHF with Passive Data Collection
The term $\big{\lVert}\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\big{\rVert}_{\mathcal{H}_{T+1}^{-1}}$ is referred to “concentrability coefficient” in the literature. It measures the distribution shift between the optimal policy and the collected data.
Remark 3.
Our result improves upon Zhu et al. (2023) in both statistical and computational efficiency. Specifically, since $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Theorem 2 directly implies a $\widetilde{\mathcal{O}}(\sqrt{d\kappa}\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ guarantee, improving their $\widetilde{\mathcal{O}}(\sqrt{d}\kappa\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ result by a factor of $\sqrt{\kappa}$. Regarding computational efficiency, their algorithm has a total storage complexity of $\mathcal{O}(T)$ and a time complexity of $\mathcal{O}(T\log T)$, leading to an amortized per-iteration cost of $\mathcal{O}(\log T)$. In contrast, our algorithm maintains a strict $\mathcal{O}(1)$ complexity per iteration, offering a substantial computational advantage.
Algorithm 3 Active Data Collection
0:  Regularization $\lambda$, step size $\eta$
1:  Initialize $\widetilde{\theta}_{1}=\bm{0}$ and $\mathcal{H}_{1}=\lambda I$
@@
3:     Choose $(x_{t},a_{t},a_{t}^{\prime})$ as Eq. (7), observe $y_{t}$
[…]
5:  end for
6:  Set $\widetilde{r}_{T+1}(x,a)=\frac{1}{T+1}\sum_{t=1}^{T+1}\phi(x,a)^{\top}\widetilde{\theta}_{t}$
6:  $\pi_{T+1}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T+1}(x,a)$
Algorithm 4 Deployment-Time Adaptation
0:  Regularization $\lambda$, step size $\eta$
1:  Initialize $\widetilde{\theta}_{1}=0$ and $\mathcal{H}_{1}=\lambda I$.
2:  for $t=1,2,\ldots,T$ do
3:     Observes the context $x_{t}$.
4:     Selects $a_{t}$ and $a_{t}^{\prime}$ as Eq. (9) and Eq. (10)
5:     Observe the preference feedback $y_{t}$
[…]
@@
5.2 Online RLHF with Active Data Collection
As established in Theorem 1, the sub-optimality gap depends on the concentrability coefficient, which quantifies the distributional mismatch between the optimal policy and the collected data. In this subsection, we propose an active data collection method that removes this dependency.
Active Data Collection.  At each iteration, we obtain a new preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$. We then update the reward model using our one-pass reward modeling method as in Eq. (5). To guide data acquisition, we actively select the query with the highest uncertainty under the current reward model. Specifically, we select the next query by solving: